Two parallel plates at a distance of 0.008m apart are maintained at a potential difference of 600v with the negative plate earthed, calculate the electric field strength around the plates.


1
Expert's answer
2021-11-22T10:13:20-0500

The electric field strength around the plates can be found as follows:


"E=\dfrac{V}{d}=\dfrac{600\ V}{0.008\ m}=7.5\times10^4\ \dfrac{V}{m}."